A person's hand pushed down with a force of 40.0N on a 8.50kg book that is sitting on a table. what is the magnitude of the normal force on the book? what is the magnitude and direction (in words) of the force on the person's hand due to the book ?
1
Expert's answer
2019-10-31T11:15:01-0400

The magnitude of the normal force on the book:


"N=F+mg=40+(8.5)(9.8)=123\ N"

The magnitude and direction of the force on the person's hand due to the book: 40.0 N upwards.